OpenStudy (anonymous):

A baseball is thrown upward from ground level with an initial velocity of 48 feet per second, and its height h, in feet, is given by: h(t)=-16t^2+48t where t is time in seconds: Determine how many seconds after the baseball is thrown it reaches its maximum height. Then find the maximum height the baseball can reach. Need help from experienced tutors, thankyou

OpenStudy (anonymous):

the vertex of any quadratic \(y=ax^2+bx+c\) has the first coordinate \(-\frac{b}{2a}\) in your case \(a=-16,b=48\) and so \(-\frac{b}{2a}=-\frac{48}{2\times (-16)}=\frac{48}{32}=\frac{3}{2}\)

OpenStudy (anonymous):

so one and a half seconds since you know the time, you know the height by replacing \(t\) by \(\frac{3}{2}\) and see what you get

OpenStudy (anonymous):

how would you find t

OpenStudy (anonymous):

by plugging it in....right?

OpenStudy (anonymous):

ok i got it thankyou